The time for a human pregnancy is distributed

N(39, 1.542)

weeks in the US.

a) A preterm baby is defined as a baby born earlier than 37 weeks. What percentage of babies are preterm?
Use 4 decimals.


b) A postmature baby is born after more than 41 weeks. What percentage of babies are postmature?
Use 4 decimals.


c) What percentage are neither preterm nor postmature?
Use 4 decimals.


d) My wife says those women who have the longest 3% of pregnancies must be pretty miserable. How long do their pregnancies last?
Use 2 decimals.
___________ weeks.

e) My first son was born early. In fact only 39% of babies are born earlier than he was. How many weeks was it before he was born?
Use 2 decimals.
___________weeks.
Answer #1

a)

for normal distribution z score =(X-μ)/σx
here mean=       μ= 39
std deviation   =σ= 1.5420
probability = P(X<37) = P(Z<-1.3)= 0.0968

b)

probability = P(X>41) = P(Z>1.30)= 1-P(Z<1.3)= 1-0.9032= 0.0968

c)

percentage are neither preterm nor postmature =1-0.0968-0.0968=0.8064

d)

for 97th percentile critical value of z= 1.88
therefore corresponding value=mean+z*std deviation= 41.90

e)

for 39th percentile critical value of z= -0.28
therefore corresponding value=mean+z*std deviation= 38.57